# Debt Behind the AI Boom: A Large-Scale Empirical Study of AI-Generated Code in the Wild

### 30 Mar 2026 v1

@@ 5.2. RQ2: Comparison Across AI Coding Assistants
Figure 7. Net impact of AI coding assistants: issues introduced vs. fixed by issue type.
5.3. RQ3: Lifecycle of AI-Introduced Debt
@@
Net Impact.  In RQ1 and RQ2, we focus on the technical debt introduced by AI-authored commits. However, AI coding assistants can also remove existing issues during refactoring or code improvement. To better understand the overall lifecycle of AI-introduced debt, we compare the number of issues introduced and fixed by AI commits (see Figure 7). For code smells, we can see that AI-authored commits fix more issues than they introduce (449,984 vs. 431,850), resulting in a net reduction of 18,134 code smells. In contrast, for runtime bugs and security issues, AI commits introduce more issues than they fix. What is interesting is that AI introduces nearly twice as many security issues as it fixes. These findings indicate that the net impact of AI coding assistants is mixed. AI coding assistants can help reduce maintainability issues, which tend to follow simple and repetitive patterns. However, for bugs and security issues, which require a deeper understanding and reasoning about program logic and context, AI coding assistants introduce more problems than they resolve.
@@
Figure 8. Cumulative growth of AI-introduced issues over time, by issue type.
Table 6. Issue survival by time since introduction. Values show survival rates and the number of surviving issues per 100 AI-authored commits.
@@
Time Since Introduction Survival Rate Surviving Issues per 100 Commits
Overall Code Smells Bugs Security
>> 9 months 19.20% 22.20 19.58 0.95 1.67
6–9 months 17.90% 29.73 25.60 1.56 2.57
3–6 months 29.30% 41.23 35.62 2.62 2.99
<< 3 months 24.80% 39.92 33.03 3.10 3.79
All Issues 24.20% 37.25 31.49 2.55 3.22
@@
Issue Survival.  The net impact analysis above provides an overview of what AI coding assistants add and remove. But it does not show what happens to the specific issues introduced by AI. To answer this question, we track each AI-introduced issue to the latest repository snapshot and check whether it still exists at HEAD. Figure 8 shows that the cumulative number of surviving issues keeps growing over time. The total volume of unresolved technical debt increases rapidly, climbing from just a few hundred issues in early 2025 to over 110,000 surviving issues by February 2026. This suggests that as the rapid adoption of AI coding assistants continues, the amount of AI-introduced debt in real-world repositories is also growing significantly.
Table 6 provides a normalized view of this surviving debt. Overall, we can see that 24.2% of tracked AI-introduced issues potentially still survive at HEAD, which corresponds to 37.25 surviving issues per 100 AI-authored commits. Survival rates also differ across issue types. Security issues are the most likely to remain at HEAD (41.1%), followed by runtime bugs (30.3%) and code smells (22.7%). When we group these issues by age, we can see an interesting pattern. It is apparent that issues introduced earlier tend to have fewer survivors per 100 commits than more recent ones. For issues older than 9 months, only 22.20 issues per 100 commits still survive, compared to 39.92 for those introduced in the last 3 months. A similar decrease can also be observed for all three issue types. This suggests that AI-introduced debt does get gradually resolved over time. However, the resolution process is slow, and even after nine months, a substantial amount of debt remains in the codebase.
@@
Figure 9. A TypeScript lint issue introduced by a Claude-authored commit in Stirling-PDF was fixed one day later by removing the unused variable.
@@
These aggregate results are also reflected in real-world repositories. For example, the broad exception handling issue in Listing 1 was fixed within hours after it was introduced (archivebox, 2025c, b). Similarly, in Stirling-PDF (Tools, 2026) (>>75k stars), a Claude-authored commit introduced an unused variable filename in a TypeScript file (Tools, 2025b). As shown in Figure 9, the maintainer fixed it the next day with a commit titled “Fix TypeScript linting error” (Tools, 2025a). In contrast, the undefined variable bug in firecrawl (Listing 2) took 42 days before the maintainer fixed it (firecrawl, 2025a). However, some issues can survive for much longer or even remain unresolved. For instance, a Devin-authored commit in brave_search_tool.py added a call to requests.get(...) without a timeout on December 2024 (crewAIInc, 2024). This is a known potential security issue, because requests made without a timeout may block indefinitely if the remote service does not respond(PyCQA, 2023). However, the issue still remains in the latest repository revision.
Answer to RQ3: AI-introduced debt is not always removed quickly after introduction. Overall, 24.2% of tracked issues still survive at HEAD, and even issues older than 9 months leave a substantial surviving burden in the codebase.
6. Discussion
6.1. Implications
@@
The main implication is not just that AI coding assistants may produce low-quality code. More importantly, AI-assisted software development changes how technical debt enters and remains in production systems. Prior studies have shown that developers are more likely to over-trust AI suggestions (Sabouri et al., 2025). This over-trust can lead to a higher acceptance rate of AI-generated code, even when it contains issues. It means that many AI-introduced issues can accumulate in the codebase. Our results show that code smells are the most common type of AI-introduced debt. They often do not break the software system immediately, making them easy to accept during code review. But AI coding assistants allow developers to produce code at much higher speed and volume. As a result, these minor issues can accumulate into a substantial maintenance burden over time. Figure 8 shows that the cumulative number of surviving AI-introduced issues continues to rise over time, exceeding 110,000 by February 2026. The technical debt introduced by AI does not have to be a temporary side effect. It can become part of long-term maintenance challenges for modern software systems.
@@
At the same time, our findings suggest although AI coding assistants introduce technical debt, they also fix existing issues in the codebase. First, we observe that AI co-authored commits actually fix a similar number of code smell issues as they introduce. This suggests that AI coding assistants are able to perform local cleanup and repetitive maintenance tasks effectively. They can recognize and address surface-level code quality problems (e.g., formatting, naming, or simple refactoring opportunities). However, what is concerning is that AI coding assistants seem to be less effective at fixing bugs and security issues, and they even introduce more of these than they fix. Section 5.3 shows that the survival rate of AI-introduced runtime bugs and security issues is much higher than that of code smells. This inconsistency suggests that AI coding assistants may struggle with changes that require deeper understanding of program behavior, execution context, or security implications. Also, the practical impact of bugs and security issues is often more severe than that of code smells, making it more critical to address them effectively. Developers should not treat all AI-generated code as equally trustworthy, and they should pay particular attention to changes that may introduce bugs or security vulnerabilities.
@@
Our cross-tool comparison shows that this problem cannot be solved simply by switching from one assistant to another. Figure 6 shows that all five tools introduce a similar pattern of issues. They all have a high rate of code smells, and a non-trivial rate of bugs and security issues. This suggests that the quality risk is a systemic issue with the current mode of AI-assisted development. Thus, developers and teams should carefully review AI-generated code regardless of the tool used. Static analysis, tests, and security checks should be part of the normal workflow. Code review should extend beyond the point of merge. The main reason is that our study shows that 24.2% of tracked AI-introduced issues still survive at HEAD, and older issues are not fully cleaned up even after months. Merging the AI-generated code does not mean the end of the story since debt can persist and accumulate over time. This makes continuous monitoring and targeted debt repayment necessary for AI-touched code.
@@
Overall, our results suggest that future research and tool design should not just focus on generating more acceptable code. We also need to ask whether that code remains maintainable, correct, and secure over time. Most existing research studies (Liu et al., 2024; Siddiq et al., 2024) focus on short-term outcomes such as task completion, acceptance rate, or immediate correctness. However, these measures only capture what happens when the code is introduced, not what happens later in maintenance. Future work needs to examine which factors make AI-introduced debt more likely to persist (e.g., repository maturity, review intensity, or task type). Our findings also suggest the need for better assistants. Future tools should make stronger checks for security-sensitive changes, use repository context more effectively, and clearly show AI provenance so reviewers can better judge the risk of a change. In the end, the key question is not only whether AI can produce code at scale, but whether the software engineering ecosystem can manage that code well over time.
6.2. Limitations
Lack of human-written baseline.  Our study does not compare AI-authored commits against a clean baseline of human-written commits. In real world, developers may use AI coding assistants but they do not leave any trace of it in the Git metadata. Even when a commit is labeled as AI-authored, the final code may include contributions from the human developer (e.g. edits, additions). This makes it difficult to construct a reliable baseline of purely human-authored commits. Further, comparing our datasets with an unreliable baseline would cause biased results and misleading conclusions. For this reason, we focus on tracking the technical debt inside explicitly confirmed AI-authored or co-authored commits.
Scope of technical debt.  In this study, we focus on code-level technical debt that can be detected by static analysis tools. We categorize issues into three types: code smells, bugs, and security issues. This categorization is based on the capabilities of the tools we use (e.g., Pylint, Bandit) and the common types of technical debt in software development. However, technical debt is a broader concept that includes architectural debt, design debt, documentation debt, and other forms of maintenance challenges. Thus, our findings should be interpreted as reflecting the code-level technical debt introduced by AI coding assistants, rather than the full spectrum of technical debt that may arise in software projects.
Survival Tracking Limitations.  Our issue survival analysis determines whether an issue still exists at HEAD by matching rule identifiers and surrounding code context. However, a file may be deleted or entirely rewritten between the introducing commit and HEAD. In such cases, the issue is classified as resolved, even though the resolution may not be a deliberate fix. We acknowledge this limitation and note that our survival rates may slightly underestimate the true persistence of AI-introduced debt.
7. Related work
@@
Code Quality of AI-Generated Code.  Previous studies have examined the quality and security of AI-generated code. Based on controlled experiments, they showed that AI coding assistants (e.g., GitHub Copilot and ChatGPT) are able to produce functional code, but the code quality varies widely across languages, tasks, and prompts (Nguyen and Nadi, 2022; Liu et al., 2024; Mastropaolo et al., 2023b). AI-generated code can also contain security weaknesses, and developers may over-trust it and fail to properly review it (Pearce et al., 2025; Perry et al., 2023; Sandoval et al., 2023). Recent studies have also begun examining AI-generated code in real-world production environments. They show that AI-generated code is being widely adopted in platforms such as GitHub and Stack Overflow, and that it can carry quality and security issues (Fu et al., 2025; Siddiq et al., 2024; Wang et al., 2025). He et al. (He et al., 2025) studied the impact of Cursor adoption on 807 repositories and observed persistent increases in code complexity. Watanabe et al. (Watanabe et al., 2025) found that most Claude Code pull requests are merged, though many require human revisions. However, these studies mainly focus on a single tool or a narrow set of quality issues, and they do not track how those issues evolve.
@@
Technical Debt in Software Development.  Technical debt is a core topic in software engineering research for years. Previous work has introduced many automated methods to identify code smells and self-admitted technical debt in software repositories (Yan et al., 2018; Ren et al., 2019; Palomba et al., 2018). Beyond detection, many studies have also investigated the lifecycle of technical debt in human-written code. Tufano et al. (Tufano et al., 2017) showed that code smells are often introduced during normal development. Once introduced, they can linger in the codebase for a long time before anyone removes them (Tufano et al., 2017). Digkas et al. (Digkas et al., 2018) identified a similar pattern in large open-source ecosystems. Their results showed that technical debt mainly accumulates when new code is added (Digkas et al., 2018, 2020). Other studies further suggest that debt is rarely removed in an intentional way, and that automated repayment is still difficult in practice (Zampetti et al., 2018; Mastropaolo et al., 2023a). Together, these studies provide a strong foundation for understanding technical debt in human-written software.
8. Threats to Validity
@@
External validity. Our study focuses on public GitHub repositories with at least 100 stars and production source files in Python, JavaScript, and TypeScript. Therefore, our findings may not generalize to private repositories, smaller projects, or software written in other languages. In addition, we only analyze AI-authored commits that leave explicit traces in Git metadata. Projects or tools that use AI without such traces are outside the scope of our dataset.
@@
Internal validity. Our pipeline depends on the correctness of AI attribution, issue matching, and lifecycle tracking. Although we use explicit Git metadata rather than proxy signals, some commits may still include both AI and human contributions. Similarly, matching issues across revisions is challenging because files evolve over time and issues may shift location. To reduce this risk, we compare code before and after each commit, restrict introduced issues to changed lines, and use rule identifiers, messages, and surrounding code context during matching. However, some attribution errors may still remain.
@@
Construct validity. Technical debt is a broad concept with many possible forms. In this study, we operationalize technical debt mainly through code smells, correctness issues, and security issues detected by static analysis tools. This choice allows us to measure debt consistently at scale and track it over time at the commit level. However, static analysis tools can produce false positives, flagging code that is technically correct but matches a known risky pattern. Additionally, static analysis does not cover all forms of technical debt. Architectural debt, design erosion, documentation debt, and test adequacy issues are outside the scope of our tools. Our results should therefore be interpreted as evidence about code-level technical debt, not the full spectrum of quality challenges that AI-generated code may introduce.
9. Conclusion

### 26 Apr 2026 v2

@@ V-B RQ2: Comparison Across AI Coding Assistants
Figure 8: Net impact of AI coding assistants: issues introduced vs. fixed by issue type.
V-C RQ3: Persistence of AI-Introduced Debt
@@
Net Impact.  In RQ1 and RQ2, we focus on the technical debt introduced by AI-authored commits. However, AI coding assistants can also remove existing issues during refactoring or code improvement. To better understand the overall lifecycle of AI-introduced debt, we compare the number of issues introduced and fixed by AI commits (see Figure 8). For code smells, we can see that AI-authored commits fix more issues than they introduce (439,817 vs. 432,748), resulting in a net reduction of 7,069 code smells. In contrast, for correctness and security issues, AI commits introduce more issues than they fix. What is interesting is that AI introduces about 1.5 times as many security issues as it fixes. These findings indicate that the net impact of AI coding assistants is mixed. AI coding assistants can help reduce maintainability issues, which tend to follow simple and repetitive patterns. However, for correctness and security issues, which require a deeper understanding and reasoning about program logic and context, AI coding assistants introduce more problems than they resolve.
@@
Figure 9: Cumulative growth of AI-introduced issues over time, by issue type.
TABLE VI: Survival of AI-introduced issues by time since introduction.
@@
Time Since Introduction # Commits # Issues # Surviving Survival Rate
>> 9 months 17,612 21,421 4,893 22.8%
6–9 months 56,955 89,135 17,307 19.4%
3–6 months 79,986 111,551 31,492 28.2%
<< 3 months 147,979 242,793 51,672 21.3%
All Issues 302,532 464,900 105,364 22.7%
@@
Issue Survival.  The net impact analysis above provides an overview of what AI coding assistants add and remove. But it does not show what happens to the specific issues introduced by AI. To answer this question, we track each AI-introduced issue to the latest repository snapshot and check whether it still exists at HEAD. Figure 9 shows that the cumulative number of surviving issues keeps growing over time. The total volume of unresolved technical debt increases rapidly, climbing from just a few hundred issues in early 2025 to over 100k surviving issues by February 2026. This suggests that as the rapid adoption of AI coding assistants continues, the amount of AI-introduced debt in real-world repositories is also growing significantly.
Table VI provides an age-cohort view of issue survival. Overall, 105,364 out of 464,900 tracked AI-introduced issues still survive at HEAD, corresponding to a survival rate of 22.7%. Surviving issues appear in all age cohorts, including issues introduced more than nine months earlier. For example, 4,893 issues introduced more than nine months ago still remain at HEAD. The survival rate varies across cohorts, ranging from 19.4% for issues introduced 6–9 months ago to 28.2% for issues introduced 3–6 months ago. This suggests that AI-introduced debt is not always removed quickly after it enters the codebase. Although the cohort-level survival rates do not show a simple monotonic trend, the main finding is clear: a substantial number of AI-introduced issues remain unresolved over time.
@@
Figure 10: A TypeScript lint issue introduced by a Claude-authored commit in Stirling-PDF was fixed one day later by removing the unused variable.
@@
These aggregate results are also reflected in real-world repositories. For example, the broad exception handling issue in Listing 1 was fixed within hours after it was introduced [5, 4]. Similarly, in Stirling-PDF [67] (>>75k stars), a Claude-authored commit introduced an unused variable filename in a TypeScript file [66]. As shown in Figure 10, the maintainer fixed it the next day with a commit titled “Fix TypeScript linting error” [65]. In contrast, the undefined variable bug in firecrawl (Listing 2) took 42 days before the maintainer fixed it [16]. In some cases, some issues can survive for much longer or even remain unresolved. For instance, a Devin-authored commit in brave_search_tool.py added a call to requests.get(...) without a timeout in December 2024 [8]. This is a known potential security issue, because requests made without a timeout may block indefinitely if the remote service does not respond [51]. However, the issue still remains in the latest repository revision.
Answer to RQ3: AI-authored commits fix slightly more code smells than they introduce, but introduce more correctness and security issues than they fix. Overall, 22.7% of tracked AI-introduced issues still survive at HEAD, including issues introduced more than nine months earlier.
VI Discussion
VI-A Implications
@@
AI-assisted development creates persistent debt, not just temporary low-quality code. The main implication is not just that AI coding assistants may produce low-quality code. More importantly, AI-assisted software development changes how technical debt enters and remains in production systems. Prior studies have shown that developers are more likely to over-trust AI suggestions [56]. This over-trust can lead to a higher acceptance rate of AI-generated code, even when it contains issues. It means that many AI-introduced issues can accumulate in the codebase. Our results show that code smells are the most common type of AI-introduced debt. They often do not break the software system immediately, making them easy to accept during code review. But AI coding assistants allow developers to produce code at a much higher speed and volume. As a result, these minor issues can accumulate into a substantial maintenance burden over time. Figure 9 shows that the cumulative number of surviving AI-introduced issues continues to rise over time, exceeding 100k by February 2026. The technical debt introduced by AI does not have to be a temporary side effect, but a long-term maintenance challenge for modern software systems.
@@
Developers should be especially cautious about correctness and security issues. Our findings suggest that although AI coding assistants introduce technical debt, they also fix existing issues in the codebase. First, we observe that AI co-authored commits actually fix a similar number of code smell issues as they introduce. This suggests that AI coding assistants are able to perform local cleanup and repetitive maintenance tasks effectively. They can recognize and address surface-level code quality problems (e.g., formatting, naming, or simple refactoring opportunities). However, what is concerning is that AI coding assistants seem to be less effective at fixing correctness and security issues, and they even introduce more of these than they fix. Section V-C shows that many AI-introduced issues still survive at HEAD, while the net-impact analysis shows that AI commits introduce more correctness and security issues than they fix. This inconsistency suggests that AI coding assistants may struggle with changes that require a deeper understanding of program behavior, execution context, or security implications. Also, the practical impact of correctness and security issues is often more severe than that of code smells, making it more critical to address them effectively. Developers should not treat all AI-generated code as equally trustworthy, and they should pay particular attention to changes that may introduce correctness or security vulnerabilities.
@@
Technical debt cannot be solved by switching between AI coding tools. Our cross-tool comparison shows that this problem cannot be solved simply by switching from one assistant to another. Figure 7 shows that all five tools introduce a similar pattern of issues. They all have a high rate of code smells, and a non-trivial rate of correctness and security issues. This suggests that the quality risk is a systemic issue with the current mode of AI-assisted development. Thus, developers and teams should carefully review AI-generated code regardless of the tool used. Static analysis, tests, and security checks should be part of the normal workflow. Code review should extend beyond the point of merge. The main reason is that our study shows that 22.7% of tracked AI-introduced issues still survive at HEAD, and older issues are not fully cleaned up even after months. Merging the AI-generated code does not mean the end of the story since debt can persist and accumulate over time. This makes continuous monitoring and targeted debt repayment necessary for AI-touched code.
@@
Future research and tool design should target long-term code health. Our results suggest that future research and tool design should not just focus on generating more acceptable code. We also need to ask whether that code remains maintainable, correct, and secure over time. Most existing research studies [35, 62] focus on short-term outcomes such as task completion, acceptance rate, or immediate correctness. However, these measures only capture what happens when the code is introduced, not what happens later in maintenance. Future work needs to examine which factors make AI-introduced debt more likely to persist (e.g., repository maturity, review intensity, or task type). Our findings also suggest the need for better assistants. Future tools should make stronger checks for security-sensitive changes, use repository context more effectively, and clearly show AI provenance so reviewers can better judge the risk of a change. In the end, the key question is not only whether AI can produce code at scale, but whether the software engineering ecosystem can manage that code well over time.
VII Related work
@@
Code Quality of AI-Generated Code.  Previous studies have examined the quality and security of AI-generated code. Based on controlled experiments, they showed that AI coding assistants (e.g., GitHub Copilot and ChatGPT) are able to produce functional code, but the code quality varies widely across languages, tasks, and prompts [44, 35, 38]. AI-generated code can also contain security weaknesses, and developers may over-trust it and fail to properly review it [47, 49, 57]. Recent studies have also begun examining AI-generated code in real-world production environments. They show that AI-generated code is being widely adopted in platforms such as GitHub and Stack Overflow, and that it can carry quality and security issues [20, 62, 69]. He et al. [28] studied the impact of Cursor adoption on 807 repositories and observed persistent increases in code complexity. Watanabe et al. [70] found that most Claude Code pull requests are merged, though many require human revisions.
Together, these studies show that AI-generated code can contain quality and security problems in both controlled and real-world settings. However, these studies mainly focus on a single tool or a narrow set of quality issues, and they do not track how those issues evolve. In contrast, our work provides a comprehensive analysis of the quality and security of AI-generated code across multiple tools, and tracks how those issues persist in production repositories over time.
Empirical Studies of AI-Assisted Development Practices.  Recent studies have examined how AI coding assistants are used in real software development practice. Several studies examined the productivity impact of these tools. Peng et al. [48] conducted a randomized controlled trial and found that developers using GitHub Copilot completed tasks 55.8% faster. At industrial scale, Ziegler et al. [73] and Murali et al. [42] reported that suggestion acceptance rate strongly correlates with self-reported productivity, with acceptance rates around 22–30%. Other studies focus on usage patterns and developer perceptions. Liang et al. [34] surveyed 410 developers and found that developers mainly use AI assistants to reduce keystrokes and recall syntax, but often reject suggestions that fail to meet functional requirements [12]. Sabouri et al. [56] further observed that developers keep only 52% of AI suggestions after review, and Klemmer et al. [32] found that developers widely use AI assistants for security-critical tasks despite concerns about suggestion quality.
These studies improve our understanding of how developers use AI assistants in practice. However, they mainly focus on adoption, usability, trust, productivity, and collaboration, rather than on the technical debt carried by AI-authored code. In contrast, our work examines the code-level debt introduced by AI coding assistants and studies whether that debt persists in production repositories over time.
@@
Technical Debt in Software Development.  Technical debt has been a core topic in software engineering research for years. Previous work has introduced many automated methods to identify code smells and self-admitted technical debt in software repositories [71, 54, 46]. Beyond detection, many studies have also investigated the lifecycle of technical debt in human-written code. Tufano et al. [68] showed that code smells are often introduced during normal development. Once introduced, they can linger in the codebase for a long time before anyone removes them [68]. Digkas et al. [14] identified a similar pattern in large open-source ecosystems. Their results showed that technical debt mainly accumulates when new code is added [14, 13]. Other studies further suggest that debt is rarely removed in an intentional way, and that automated repayment is still difficult in practice [72, 37].
Prior studies provide a strong foundation for understanding technical debt in human-written software. However, it remains unclear whether AI-generated code follows the same patterns, and whether the same tools and techniques can be applied to it. Our work fills this gap by analyzing the technical debt in AI-generated code and exploring how it persists in production repositories over time.
VIII Threats to Validity
@@
External Validity. Our study focuses on public GitHub repositories with at least 100 stars and production source files in Python, JavaScript, and TypeScript. Therefore, our findings may not generalize to private repositories, smaller projects, or software written in other languages. In addition, we only analyze AI-authored commits that leave explicit traces in Git metadata (e.g., bot actor logins, AI author emails, or Co-authored-by trailers). AI-assisted contributions that leave no such trace are outside the scope of our dataset. This design is analogous in spirit to research on self-admitted technical debt (SATD), which studies the subset of technical debt that developers explicitly document rather than the full universe of debt [50, 11]. Like SATD research, our findings characterize a visible and attributable subset rather than the entire population of AI-assisted work. We also do not compare AI-authored commits against a baseline of purely human-written commits. Because developers may use AI without leaving any Git trace, and even AI-labeled commits may mix human edits, a reliable human-only baseline is difficult to construct, and comparing against an unreliable baseline could bias the results. We therefore focus on tracking technical debt inside explicitly confirmed AI-authored or co-authored commits.
@@
Internal Validity. Our pipeline depends on the correctness of AI attribution, issue matching, and lifecycle tracking. Although we use explicit Git metadata rather than proxy signals, some commits may still include both AI and human contributions. Similarly, matching issues across revisions is challenging because files evolve over time and issues may shift location. To reduce this risk, we compare code before and after each commit, restrict introduced issues to changed lines, and use rule identifiers, messages, and surrounding code context during matching. For the issue survival analysis, a file may be deleted or entirely rewritten between the introducing commit and HEAD. In such cases, the issue is classified as resolved, even though the resolution may not be a deliberate fix. Our reported survival rates may therefore slightly underestimate the true persistence of AI-introduced debt.
@@
Construct Validity. Technical debt is a broad concept with many possible forms. In this study, we operationalize technical debt mainly through code smells, correctness issues, and security issues detected by static analysis tools. This choice allows us to measure debt consistently at scale and track it over time at the commit level. However, static analysis tools can produce false positives, flagging code that is technically correct but matches a known risky pattern. In addition, our security findings include both active vulnerabilities and latent unsafe patterns (i.e., security debt). We do not separate these quantitatively, but Section V-A illustrates both cases. Static analysis also does not cover all forms of technical debt. Architectural debt, design erosion, documentation debt, and test adequacy issues are outside the scope of our tools. Our results should therefore be interpreted as evidence about code-level technical debt, not the full spectrum of quality challenges that AI-generated code may introduce.
IX Conclusion and Future Work